# Scaling-aware rating of count forecasts

### v3

@@ 2.4.2 Ranked Probability Score
MAE has an immediate business interpretation: A planner who orders $\text{median}(r_{j})$ pieces will, on average, have an excess or understock of MAE pieces. The necessity to use the median of the distribution, however, leads to unpleasant discontinuous behavior (the median of the Poisson distribution jumps from 0 to 1 at the rate $r=\log 2\approx 0.693$). To provide a business interpretation while acknowledging the probabilistic nature of the forecast, the forecaster also evaluates the Mean Ranked Probability Score (MRPS), which is the mean over all product-location-days of the discrete Ranked Probability Score (RPS) (Epstein, 1969; Gneiting and Raftery, 2007; Tichy, 2022).
@@ 2.6 Baseline model and naïve scaling trap
To obtain a maximally sharp and calibrated baseline model, we construct predictions $r^{\text{baseline}}_{j}$ via in-sample Expectation Maximization (EM) of the Poisson likelihood, applied on the entire set of sales $\vec{s}$ (see A). That is, the EM procedure ingests all outcomes $s_{j}$; using the empirical frequencies $P(s)$ and the Poisson distribution assumption (4), it computes a prior distribution of rates $P(r)$, from which it samples the rate $r^{\text{baseline}}_{j}$ for each outcome $s_{j}$ via $P(r|s)=P(s|r)P(r)/P(s)$. The rate $r_{j}$ could hence have generated the outcome $s_{j}$ under a Poisson distribution. The Expectation Maximization procedure does not use any feature or any structure in the data (e.g., whether two sales values $s_{j}$, $s_{k}$ belong to the same product in different stores or to the same product-store-combination for different days is not accounted for). In summary, we obtain a baseline “forecast” (given that it is generated and evaluated using the same dataset, it is not, technically speaking, a forecast of some future event) $\vec{r}^{\text{baseline}}$ for which the sales $s_{j}$ are Poisson distributed for each prediction $r^{\text{baseline}}_{j}$.
The calibration diagrams of Figure 1 confirm that the baseline $\vec{r}^{\text{baseline}}$ is an excellent model, since the unconditioned and the conditioned distributions of actually observed sales $s$ match the Poisson prediction well. It is fair to say that this baseline forecast is the best that one could possibly expect under the given set of product-related covariates Tichy (2023c). We will see in Section 5 that the state-of-the-art in commercial machine learning indeed comes close to the baseline performance.
@@
A direct interpretation of the forecast metrics across departments (or any other way of segregating the data) is deeply problematic and is doomed to misalign efforts for forecast improvement. Since the “effect” that one sees is infected by precision scaling, we propose to designate such distorsion as the naïve scaling trap.
@@
3 Rating metrics in a scaling-aware way
@@ 3.1 Revealing and avoiding the naïve scaling trap
If all predictions are aggregated globally, the resulting expectation value of the metric will be mostly borne by fast-movers. Therefore, an additional segregation by selling rate is useful: We group the pairs of predictions $r_{j}$ and observations $s_{j}$ into buckets, the bucket of a pair is chosen according to the prediction value $r_{j}$. If one chose the outcome $s_{j}$ to define the bucket, one would suffer from a hindsight selection bias (Tichy and Feindt, 2022), which leads to the forecaster’s dilemma (Lerch et al., 2017; Tichy, 2023b, 2024). Logarithmically spaced buckets are useful in retail to avoid cases in which one bucket contains the majority of all predictions, i.e., we group by
$R_{j}=\frac{1}{n_{\text{bins}}}\text{round}\left(\text{log}_{10^{\left(\frac{1}{n_{\text{bins}}}\right)}}r_{j}\right),$ (10)
where we round to integers such that there are $n_{\text{bins}}$ bins between two powers of 10. Buckets are referred to by their common rounded logarithmic prediction value $R$. In practice, we have found that $n_{\text{bins}}=2$ constitutes the absolute minimum of buckets, and chosing $n_{\text{bins}}=5$ is typically a reasonable choice. The computed ratings weakly depends on the number of buckets, such that it should be chosen as a constant within one analysis.
The predictions $r_{j}$ that belong to the same bucket, indexed by $R$, come with similar expectation values of any metric. Ideally, when the prediction is truly Poissonian, we expect the achieved average metric in a bucket to match the metric’s computed expectation value. The number of bins $n_{\text{bins}}$ should thus solve the compromise between having sufficiently many predictions and observations in one bucket so we can expect measurement and expectation value to match approximately, while ensuring that all predictions in the bucket behave approximately equally. Figure 2 shows the result of this procedure for the baseline model as circles and $n_{\text{bins}}=4$, and superimposes the expectation value of the metrics under the Poisson distribution as solid line. The circle sizes reflect the square-root of the number of measurements in order to make small buckets still visible. The positions of the circles match the solid lines, confirming the Poissonian nature of our baseline forecast from Figure 1. Figure 3 differentiates further by retail department for the NMRPS: The differences in the metric values exhibited in Table 1 are driven by the differences in the distribution of rates, not by a different performance. For a given rate bucket $R$, the performance is the same for all departments, but the different buckets have different populations. Skepticism that calls into question the equal model performance in the different departments, which might have been induced by the different metric values in Table 1, can now be clearly dismissed.
@@
The intricate shapes that metric values exhibit as a function of the predicted rate in Figure 2 make comparisons between strata of data deeply problematic. The “best case” value of a metric depends strongly on the prediction itself. Faster-selling departments naturally achieve higher values of MAE, MRPS and lower values of WMAPE, NMRPS, as exemplified by Figure 3. Statements like “we achieve 38% WMAPE on Foods 3, but 78% WMAPE on Hobbies 2, we need to improve Hobbies 2” are almost always infected by the naïve scaling trap and need to be dismissed.
@@ 3.2 Necessity for imperfect reference forecasts
The setting so far was artificial: The baseline model is ideal, and unsurprisingly matches the theoretical expectation; we merely confirmed that the Expectation Maximization algorithm works. Figure 3 explains why model performance across departments is only ostensibly different. Characterizing the ideal case is, however, not sufficient for judging forecasts in practice: One needs to quantify how “close” the model performance is to ideal, and rate this distance to conclude whether the departure from the ideal case is acceptable or problematic. For this purpose, we need to define imperfect benchmarks that set the standard for “excellent”, “good”, “OK”, “fair”, ”insufficient” and “unacceptable” performance, both regarding the level of bias and the level of noise.
@@
There are, however, many ways to be wrong: A forecast can be biased, it can be affected by noise, suffer from both, or by more complicated artifacts (e.g., finite-stock effects, zero-inflation…). No general “worst benchmark model” can be found for unbounded metrics that we are dealing with: Given a forecast, one can construct arbitrarily incompatible (“bad”) observations; given a set of observations, one can construct arbitrarily terrible forecasts. A universal definition of “excellent” to “unacceptable” models that would fit all industries is unlikely to be ever possible (Kolassa, 2008). We therefore expose the industry-specificity by few parameters within a simple model for imperfect reference forecasts. As a “bad” reference forecast, we use a global forecast which does not individualize at all, but produces the same mean value for all products, locations and days.
@@ 3.3 Scaling of imperfect reference forecasts
Given predictions of $r_{j}\approx 10$ ($R=1$), we ideally expect the histogram of observations to match the Poisson distribution. To simplify judging the deviation from that ideal, we use negative-binomial distributions of different variances to set reference ratings, examples of which are shown in Figure 4. We assign the variances at prediction $r_{j}=10$ and the overall bias factors (mean prediction / mean observation) a qualitative score:
@@
The ratings for bias and for variance are independent, that is, a forecast can enjoy Poisson-like variance, but be quite biased, or vice versa. The 12 parameters fully define our rating, they were chosen such that a naïve model for the M5 competition data is rated “fair” by most metrics.
Figure 4: Overdispersed distributions that define the reference forecasts “perfect” (Poissonian, blue), “excellent” (light green), “good” (dark green), etc., for a predicted rate $\mu=10$. When the forecaster predicts a Poisson-distributed target, we rate their performance with respect to a metric by referring to the metric values achieved by the respective negative binomial distributions shown in the figure.
@@
where $f$ quantifies the strength of the noise. Setting $\gamma=2$ corresponds to quadratic overdispersion (a Negative Binomial-2-process), and $\gamma=1$ to linear overdispersion (a Negative Binomial-1-process) (Hilbe, 2012; Cameron and Trivedi, 2005). In general, the variance $\sigma_{\mu}^{2}$ of the underlying distribution of the rate $\mu$, here assumed to be a $\Gamma$-distribution, is inherited by the variance of the resulting count process via $\text{variance}=\mu+\sigma_{\mu}^{2}$ (Karlis and Xekalaki, 2005). The upper right panel of Figure 5 shows the NMRPS as a function of the predicted rate, for different exponents $\gamma=1,1.5,2$. Quadratic $\gamma=2$ is unrealistic as a candidate for “equal quality”: It is too benevolent for large rates and becomes too strict for small rates. $\gamma=1$ shows the opposite behavior: It judges large rates strictly and small rates too benevolently. The intermediate value $\gamma=1.5$ provides a reasonable compromise, corroborated by this argument: When summing negative binomial-distributed random numbers, $\gamma=1$ corresponds to uncorrelated noise, $\gamma=2$ to perfectly correlated noise, and $\gamma=1.5$ partial correlation, which is realistic in the retail setting: The number of visitors of a supermarket has some uncertainty that leads to correlated, product-independent noise on the product sales, but also each individual product suffers from uncorrelated, product-dependent noise.
Figure 5: Upper left: Poisson (blue) and negative binomial (red, $\text{variance}=30$) distributions for rate $\mu=10$. Upper right: expected NMRPS as a function of rate, when the variance equals 30 at prediction 10 (corresponding to “good” as defined in Figure 4), for overdispersion exponents $\gamma=1$ (green), $\gamma=1.5$ (orange), $\gamma=2$ (black). The blue line delimits the ideally achievable (Poissonian) baseline value under overdispersion 0. Lower left: rate-10-distribution scaled down to rate 1, using linear (green), quadratic (black) and exponent-1.5 (orange) dispersion. Lower right: rate-10-distribution scaled up to rate 100, using linear, quadratic and exponent-1.5-dispersion.
@@
A given metric value, e.g., WMAPE=0.6, can be considered “perfect” for a predicted rate of 2, or “unacceptable” for a predicted rate of 100. By applying the method to different sets of data from different industries, we have experienced that large buckets found outside the “insufficient” region are indeed rare and point to severe model or data quality issues.
@@ 3.4 Summarizing bucket-wise to overall ratings
It is desirable to summarize the information contained in a set of buckets by a single number. There are two possibilities to perform this summary: One can aggregate the ratings per bucket via a mean of the bucket-wise scores $S^{(R)}$ (weighted by the number of observations per bucket), which yields an overall score $S_{\text{overall}}$ between 0% (all buckets are “unacceptable”) and 100% (all buckets are “perfect”). Alternatively, one can judge the overall aggregated metric, e.g. the overall achieved WMAPE, by setting it into the context of the metric that would have been achieved if the model quality were “excellent” to “good”. In other words, one can compute the expectation value of the metric under reference models of imperfect quality.
@@ 3.5 Ostensibly better-than-Poisson measurements
One will sometimes find buckets $R$ that achieve a “better-than-Poisson” metric, $M^{~{}(R)}<M^{~{}(R)}_{\text{perfect}}$. This surprising behavior can be due to different causes:
@@
Depending on which of these non-exclusive causes are realistic in a given application, one can decide to rate buckets for which $M^{~{}(R)}<M^{~{}(R)}_{\text{perfect}}$ by some critical value to ensure that these buckets do not push the overall rating to good values, obfuscating some problems. The method of rating the overall aggregated metric is especially prone to cancellation between smaller-than-ideal and larger-than-ideal values, which would result in an overall “excellent”-looking metric. Just like negative and positive bias in slow- and fast-movers may cancel to yield and overall ostensible good bias, “better-than-Poisson” and “worse-than-Poisson” buckets may result in an ostensibly good overall noise. Therefore, we prefer to aggregate the bucket-wise-scores as a primary tool of investigation.

### v4

@@ 2.4.2 Ranked Probability Score
MAE has an immediate business interpretation: A planner who orders $\text{median}(r_{j})$ pieces will, on average, have an excess or under-stock of MAE pieces. The necessity to use the median of the distribution, however, leads to unpleasant discontinuous behavior (the median of the Poisson distribution jumps from 0 to 1 at the rate $r=\log 2\approx 0.693$). To provide a business interpretation while acknowledging the probabilistic nature of the forecast, the forecaster also evaluates the Mean Ranked Probability Score (MRPS), which is the mean over all product-location-days of the discrete Ranked Probability Score (RPS) (Epstein, 1969; Gneiting and Raftery, 2007; Tichy, 2022).
@@ 2.6 Baseline model and naïve scaling trap
To obtain a maximally sharp and calibrated baseline model, we construct predictions $r^{\text{baseline}}_{j}$ via in-sample Expectation Maximization (EM) of the Poisson likelihood, applied on the entire set of sales $\vec{s}$ (see A). That is, the EM procedure ingests all outcomes $s_{j}$; using the empirical frequencies $P(s)$ and the Poisson distribution assumption (4), it computes a prior distribution of rates $P(r)$, from which it samples the rate $r^{\text{baseline}}_{j}$ for each outcome $s_{j}$ via $P(r|s)=P(s|r)P(r)/P(s)$. The rates $r_{j}$ that are then generated could hence have led to the outcome $s_{j}$ under a Poisson distribution. By construction, we thereby obtain a “perfect” model: For every set of rates around some value $r$, the associated observations seem to be sampled from a Poisson distribution of that rate $r$.
The Expectation Maximization procedure does not use any feature or any structure in the data (e.g., whether two sales values $s_{j}$, $s_{k}$ belong to the same product in different stores or to the same product-store-combination for different days is not accounted for). For our purpose of setting a benchmark, this baseline “forecast” (given that it is generated and evaluated using the same dataset, it is not, technically speaking, a forecast of some future event) $\vec{r}^{\text{baseline}}$ is suitable, since the sales $s_{j}$ are Poisson distributed for each prediction $r^{\text{baseline}}_{j}$.
As an alternative to using true sales data and artificially generated sales rates, one could have worked on a fully synthetic dataset. In that case, the selling rates themselves would be sampled randomly, e.g. from a gamma-distribution, and the observations $s_{j}$ would be Poisson-sampled for each rate $r_{j}$. Our point is, however, to exemplify the value of our method upon real-world datasets, which is why we use the M5 dataset as a starting point and need to generate an “ideal forecast”.
The calibration diagrams of Figure 1 confirm that the baseline $\vec{r}^{\text{baseline}}$ is an excellent model, since the unconditioned and the conditioned distributions of actually observed sales $s$ match the Poisson prediction well. It is fair to say that this baseline forecast is the best that one could possibly expect under the given set of product-related covariates (Tichy, 2023c). We will see in Section 5 that the state-of-the-art in commercial machine learning indeed comes close to the baseline performance.
@@
A direct interpretation of the forecast metrics across departments (Bower, 2023) (or any other way of segregating the data) is deeply problematic and is doomed to misalign efforts for forecast improvement. Since the “effect” that one sees is infected by precision scaling, we propose to designate such distortion as the naïve scaling trap.
@@
3 Method: Rating metrics in a scaling-aware way
@@ 3.1 Revealing and avoiding the naïve scaling trap
If all predictions are aggregated globally, the resulting expectation value of the metric will be mostly borne by fast-movers. Therefore, a segregation by selling rate is useful: We group the pairs of predictions $r_{j}$ and observations $s_{j}$ into buckets, the bucket of a pair is chosen according to the prediction value $r_{j}$. If one chose the outcome $s_{j}$ to define the bucket, one would suffer from a hindsight selection bias (Tichy and Feindt, 2022), which leads to the forecaster’s dilemma (Lerch et al., 2017; Tichy, 2023b, 2024). Logarithmically spaced buckets are useful in retail to avoid cases in which one bucket contains the majority of all predictions, i.e., we group by
$R_{j}=\frac{1}{n_{\text{bins}}}\text{round}\left(n_{\text{bins}}\text{log}_{10}(r_{j})\right),$ (10)
where we round to integers such that there are $n_{\text{bins}}$ bins between two powers of 10. For example, for $n_{\text{bins}}=4$, the $R_{j}$ that can take the values $0,0.25,0.5,0.75,1,1.25$ etc. Buckets are referred to by their common rounded logarithmic prediction value $R$. In practice, we have found that $n_{\text{bins}}=2$ constitutes the absolute minimum of buckets, and choosing $n_{\text{bins}}=5$ is typically a reasonable choice. The computed ratings also weakly depend on the number of buckets, such that it should be chosen as a constant within one analysis.
The predictions $r_{j}$ that belong to the same bucket, indexed by $R$, come with similar expectation values of any metric. Ideally, when the prediction is truly Poissonian, we expect the achieved average metric in a bucket to match the metric’s computed expectation value. The number of bins $n_{\text{bins}}$ should thus solve the compromise between having sufficiently many predictions and observations in one bucket so we can expect measurement and expectation value to match approximately (i.e. a possible deviation to be considered significant), while ensuring that all predictions in the bucket behave approximately equally. Figure 2 shows the result of this procedure for the baseline model as circles and $n_{\text{bins}}=4$, and superimposes the expectation value of the metrics under the Poisson distribution as solid line. The circle sizes reflect the square-root of the number of measurements in order to make small buckets still visible. The positions of the circles match the solid lines, confirming the Poissonian nature of our baseline forecast from Figure 1. Figure 3 differentiates further by retail department for the NMRPS: The differences in the metric values exhibited in Table 1 are driven by the differences in the distribution of rates, not by a different performance. For a given rate bucket $R$, the performance is the same for all departments, but the different buckets have different populations. Skepticism that calls into question the equal model performance in the different departments, which might have been induced by the different metric values in Table 1, can now be clearly dismissed.
@@
The intricate shapes that metric values exhibit as a function of the predicted rate in Figure 2 make comparisons of traditional metrics between strata of data (Bower, 2023) deeply problematic. The “best case” value of a metric depends strongly on the prediction itself. Faster-selling departments naturally achieve higher values of MAE and MRPS, and lower values of WMAPE and NMRPS, as exemplified by Figure 3. Therefore, statements like “we achieve 38% WMAPE on Foods 3, but 78% WMAPE on Hobbies 2, we need to improve Hobbies 2” are almost always infected by the naïve scaling trap and need to be dismissed.
@@ 3.2 Necessity for imperfect reference forecasts
The setting so far was artificial: The baseline model is ideally Poissonian, calibrated, and unsurprisingly matches the theoretical expectation. We merely confirmed that the Expectation Maximization algorithm of A works. Figure 3 explains why model performance across departments is only ostensibly different. Characterizing the ideal case is, however, not sufficient for judging forecasts in practice: One needs to quantify how “close” the model performance is to ideal, and rate this distance to conclude whether the departure from the ideal case is acceptable or problematic. For this purpose, we need to define imperfect benchmarks that set the standard for “excellent”, “good”, “OK”, “fair”, ”insufficient” and “unacceptable” performance, both regarding the level of bias and the level of noise.
@@
There are, however, many ways to be wrong: A forecast can be biased, it can be affected by noise, suffer from both, or by more complicated artefacts (e.g., finite-stock effects, zero-inflation…). No general “worst benchmark model” can be found for unbounded metrics that we are dealing with: Given a forecast, one can construct arbitrarily incompatible (“bad”) observations; given a set of observations, one can construct arbitrarily terrible forecasts. A universal definition of “excellent” to “unacceptable” models that would fit all industries is unlikely to be ever possible (Kolassa, 2008). We therefore expose the industry-specificity by few parameters within a simple model for imperfect reference forecasts. As a “bad” reference forecast, we use a collective global forecast which does not individualize at all, but produces the same value for all products, locations and days.
@@ 3.3 Scaling of imperfect reference forecasts
Given predictions of $r_{j}\approx 10$ ($R=1$), we ideally expect the histogram of observations to match the Poisson distribution with rate 10. To simplify judging the deviation from that ideal, we use negative-binomial distributions of different variances to set reference ratings, with examples in Figure 4. That is, we assume that the forecaster produces a Poisson-forecast, but that the observations are generated by a negative-binomial process, such that they observe an over-dispersed distribution of outcomes. We assign the variances at prediction $r_{j}=10$ and the overall bias factors (mean prediction / mean observation) a qualitative score:
@@
The ratings for bias and for variance are independent, that is, a forecast can enjoy Poisson-like variance, but be quite biased, or vice versa. The 12 parameters fully define our rating, they were chosen such that a collective global forecast for the M5 competition data is rated “Insufficient” by most metrics.
Figure 4: Over-dispersed distributions that define the reference forecasts “perfect” (Poissonian, blue), “excellent” (light green), “good” (dark green), etc., for a predicted rate $\mu=10$. When the forecaster predicts a Poisson-distributed target, we rate their performance with respect to a metric by referring to the metric values achieved by the respective negative binomial distributions shown in the figure.
@@
where $f$ quantifies the strength of the noise. Setting $\gamma=2$ corresponds to quadratic over-dispersion (a Negative Binomial-2-process), and $\gamma=1$ to linear over-dispersion (a Negative Binomial-1-process) (Hilbe, 2012; Cameron and Trivedi, 2005). In general, the variance $\sigma_{\mu}^{2}$ of the underlying distribution of the rate $\mu$, here assumed to be a $\Gamma$-distribution, is inherited by the variance of the resulting count process via $\text{variance}=\mu+\sigma_{\mu}^{2}$ (Karlis and Xekalaki, 2005). The upper right panel of Figure 5 shows the NMRPS as a function of the predicted rate, for different exponents $\gamma=1,1.5,2$. Quadratic $\gamma=2$ is unrealistic as a candidate for “equal quality”: It is too benevolent for large rates and becomes too strict for small rates. $\gamma=1$ exhibits the opposite behavior: It judges large rates strictly and small rates too benevolently. The intermediate value $\gamma=1.5$ provides a reasonable compromise, corroborated by this argument: When summing negative binomial-distributed random numbers, $\gamma=1$ corresponds to uncorrelated noise, $\gamma=2$ to perfectly correlated noise, and $\gamma=1.5$ to partial correlation, which is realistic in the retail setting: The number of visitors of a supermarket has some uncertainty that leads to correlated, product-independent noise on the demand, but also each individual product suffers from uncorrelated, product-dependent noise.
Figure 5: Upper left: Poisson (blue) and negative binomial (red, $\text{variance}=30$) distributions for rate $\mu=10$. Upper right: expected NMRPS as a function of rate, when the variance equals 30 at prediction 10 (corresponding to “good” as defined in Figure 4), for over-dispersion exponents $\gamma=1$ (green), $\gamma=1.5$ (orange), $\gamma=2$ (black). The blue line delimits the ideally achievable (Poissonian) baseline value under over-dispersion 0. Lower left: rate-10-distribution scaled down to rate 1, using linear (green), quadratic (black) and exponent-1.5 (orange) dispersion. Lower right: rate-10-distribution scaled up to rate 100, using linear, quadratic and exponent-1.5-dispersion.
@@
A given metric value, e.g., WMAPE=0.6, can be considered “perfect” for a predicted rate of 2, or “unacceptable” for a predicted rate of 100. By applying the method to different sets of data from different industries, we have experienced that “large” buckets (containing many, more than 100, predictions and observations) that are found outside the “insufficient” region are indeed rare and point the user to severe model or data quality issues. The significance of “small” buckets (containing less than about 100 predictions and observations) is, however, questionable: Each metric inherits the randomness of the process, and will therefore be distributed around its expectation value. An unexpected metric value that a small bucket assumes is not necessarily significant, i.e. it can also be due to chance.
@@ 3.4 Summarizing bucket-wise to overall ratings
It is desirable to summarize the information contained in a set of buckets by a single number. There are two possibilities to perform this summary: One can aggregate the ratings per bucket via taking the mean of the bucket-wise scores $S^{(R)}$ (weighted by the number of observations per bucket), which yields an overall score $S_{\text{overall}}$ between 0% (all buckets are “unacceptable”) and 100% (all buckets are “perfect”). Alternatively, one can judge the overall aggregated metric, e.g. the overall achieved WMAPE, by setting it into the context of the metric that would have been achieved if the model quality were “excellent” to “good”. In other words, one can compute the expectation value of the metric under reference models of imperfect quality, answering the important question “what values of the metric could I possibly achieve?”, which we had put forward above in Section 1.1.
@@ 3.5 Ostensibly better-than-Poisson measurements
One will sometimes find buckets $R$ that achieve a “better-than-Poisson” metric, $M^{~{}(R)}<M^{~{}(R)}_{\text{perfect}}$. This unexpected behavior can be due to different causes:
@@
Depending on which of these non-exclusive causes are realistic in a given application, one can decide to rate buckets for which $M^{~{}(R)}<M^{~{}(R)}_{\text{perfect}}$ with some critical quality to ensure that these buckets do not push the overall rating to good values, obfuscating some problems. The method of rating the overall aggregated metric is especially prone to cancellation between smaller-than-ideal and larger-than-ideal values, which would result in an overall “excellent”-looking metric. Just like negative and positive bias in slow- and fast-movers may cancel to yield an ostensibly good global bias, “better-than-Poisson” and “worse-than-Poisson” buckets may result in an ostensibly good overall noise. Therefore, we prefer to aggregate the bucket-wise-scores as a primary tool of investigation.